# Bounds and Bugs: The Limits of Symmetry Metrics to Detect Partisan Gerrymandering

### v1

@@ Appendix C Short Burst Images
Figure 54: Short burst results for Pennsylvania’s Congressional map and 2016 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
@@
Figure 55: Short burst results for Pennsylvania’s Congressional map and 2016 US Senate election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
@@
Figure 56: Short burst results for Pennsylvania’s State Senate map and 2016 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
@@
Figure 57: Short burst results for Pennsylvania’s State Senate map and 2016 US Senate election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 58: Short burst results for Pennsylvania’s State House map and 2016 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
Figure 59: Short burst results for Pennsylvania’s State House map and 2016 US Senate election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
@@
Figure 60: Short burst results for Texas’s Congressional map and 2014 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
@@
Figure 61: Short burst results for Texas’s Congressional map and 2014 US Senate election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
@@
Figure 62: Short burst results for Texas’s State Senate map and 2014 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
@@
Figure 63: Short burst results for Texas’s State Senate map and 2014 US Senate election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
@@
Figure 64: Short burst results for Texas’s State House map and 2014 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.
@@
Figure 65: Short burst results for Texas’s State House map and 2014 US Senate election data, searching for maps with as many Repuglican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\min(m)$ to $0.16\min(m)$ for each metric $m$. The small numbers below each metric value range is the number of maps produced that had the corresponding number of districts won.

### v3

@@ Appendix B Images from Empirical Analysis
Figure 16: Empirical results for Massachusetts’s Congressional map and 2018 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
@@
Figure 17: Empirical results for Massachusetts’s Congressional map and 2018 US Senate election data, searching for maps with as many Republican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
@@
Figure 18: Empirical results for Massachusetts’s State Senate map and 2018 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
@@
Figure 19: Empirical results for Massachusetts’s State Senate map and 2018 US Senate election data, searching for maps with as many Republican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
@@
Figure 20: Empirical results for Michigan’s Congressional map and 2016 Presidential election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
@@
Figure 21: Empirical results for Michigan’s Congressional map and 2016 Presidential election data, searching for maps with as many Republican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
@@
Figure 22: Empirical results for Michigan’s State Senate map and 2016 Presidential election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
@@
Figure 23: Empirical results for Michigan’s State Senate map and 2016 Presidential election data, searching for maps with as many Republican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
@@
Figure 24: Empirical results for Michigan’s State House map and 2016 Presidential election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
@@
Figure 25: Empirical results for Michigan’s State House map and 2016 Presidential election data, searching for maps with as many Republican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figure 26: Empirical results for Oklahoma’s State Senate map and 2018 Gubenatorial election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figure 27: Empirical results for Oklahoma’s State Senate map and 2018 Gubenatorial election data, searching for maps with as many Republican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figure 28: Empirical results for Oklahoma’s State House map and 2018 Gubenatorial election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figure 29: Empirical results for Oklahoma’s State House map and 2018 Gubenatorial election data, searching for maps with as many Republican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figure 30: Empirical results for Oregon’s Congressional map and 2018 Gubenatorial election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figure 31: Empirical results for Oregon’s Congressional map and 2018 Gubenatorial election data, searching for maps with as many Republican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figure 32: Empirical results for Oregon’s State Senate map and 2018 Gubenatorial election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figure 33: Empirical results for Oregon’s State Senate map and 2018 Gubenatorial election data, searching for maps with as many Republican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figure 34: Empirical results for Oregon’s State House map and 2018 Gubenatorial election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figure 35: Empirical results for Oregon’s State House map and 2018 Gubenatorial election data, searching for maps with as many Republican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figure 36: Empirical results for Pennsylvania’s State Senate map and 2016 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figure 37: Empirical results for Pennsylvania’s State Senate map and 2016 US Senate election data, searching for maps with as many Republican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figure 38: Empirical results for Pennsylvania’s State House map and 2016 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figure 39: Empirical results for Pennsylvania’s State House map and 2016 US Senate election data, searching for maps with as many Republican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figure 40: Empirical results for Texas’s Congressional map and 2014 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figure 41: Empirical results for Texas’s Congressional map and 2014 US Senate election data, searching for maps with as many Republican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figure 42: Empirical results for Texas’s State Senate map and 2014 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figure 43: Empirical results for Texas’s State Senate map and 2014 US Senate election data, searching for maps with as many Republican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figure 44: Empirical results for Texas’s State House map and 2014 US Senate election data, searching for maps with as many Democratic-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.
Figure 45: Empirical results for Texas’s State House map and 2014 US Senate election data, searching for maps with as many Republican-won districts as possible. Horizontal axis is number of districts won, vertical axis is metric value ranges. The green region is from $0.16\inf(m)$ to $0.16\sup(m)$ for each metric $m$. The small number below each metric value range is the number of maps produced that had the corresponding number of districts won. The dot within each vertical bar is the mean value of that metric on all produced maps with the corresponding number of districts won.